Consider the cash flow series shown below. What value of C makes the inflow series equivalent
to the outflow series at an interest rate of 6% compounded annually?
Answer #1

We will calculate the present value of all inflow cash series and equate it with the present value of all outflow cash series.

PV= (1+r)

where PV = present value

FV = future value

r = rate of interest

and n = no. of years.

Using this and -

PV of cash inflows = PV of cash outflows

150 150 150 150 150 (1 0.06)1 (1 0.06)2(1 0.06)310.06)70.06)8 150 3C 10.06)9 0.06)4 (10.06)5 (10.06)6 (1 +0.06)10

14 1.51 + 133.50+ 125.94+ 99.76+ 94.11 88.78 = -+--+-+ 1.26 1.34 1.42 3C 1.79

683.600.79C + 0.75C+ 0.70C 1.68C

3.92C = 683.60-C-683.60/3.92 174.39